# Critical Iterative Denoising A Discrete Generative Model Applied to Graphs

### v1

@@ 3.1 Noising
$q_{t|1}(z|z_{1})=\alpha_{t}\delta_{z_{1}}(z)+(1-\alpha_{t})q_{0}(z),$ (6)
where $q_{0}(z)$ represents the noise distribution, $\delta_{z_{1}}(z)$ is the data distribution, and $\alpha_{t}$ is a non-decreasing scheduling parameter in $[0,1]$. The scheduler $\alpha_{t}$, which defines the noise level, depends on a continuous time parameter, which also takes values in $[0,1]$. Note that we adopt the notation from DFM, but Equation 6 is equivalent to the $q_{t\mid{1}}(z|z_{1})=\text{Cat}(z;{\bm{z}}_{1}^{T}\bar{{\bm{Q}}}_{t})$ from DDM seen in Section 2.2.
Unlike standard diffusion models, where noise distributions are computed via a Markov chain (Eq. 1), our noisy distributions are directly parameterized as a mixture of the noise and the data distribution. Thus we do not assume any time dependency in the noising process other than the dependency on the original data point $z_{1}$. This assumption introduces a significant simplification over DDM, leading to beneficial implications for the ID denoising process. In figure LABEL:fig:noising we juxtapose the graphical models of our method and DDM at the noising phase. More formally we assume conditional independence of the noisy distributions across time given $z_{1}$:
@@
The noisy distributions are conditionally independent such that:
$q_{t|1}(z|z_{1})=q_{t|1}(z|z_{s},z_{1})\quad\forall t\neq s.$ (7)
As a result, our noising procedure is not a diffusion process. Under this assumption, the denoising phase does not suffer from error propagation and accumulation caused by the temporal dependencies inherent in standard DDM (see Section 3.2).
@@
The proposition follows from the fact that our noising procedure corresponds to the close-form expression (Eq. 3) used in discrete diffusion to sample efficiently from $q_{t|1}(z|z_{1})$. However, we make no assumption that the noisy distribution results from a diffusion process.
@@ 3.2 Denoising
Denoising progressively refines noisy instances by learning the distribution $p_{s|t}(z|Z_{t})$, $s=t+\Delta t$. By leveraging Assumption 3.1, the denoising process is significantly simplified compared to discrete diffusion, as formalized in the following proposition:
@@
We can factorize the denoising process as:
$p_{s|t}(z|Z_{t})=\sum_{z_{1}}q_{s|1}(z|z_{1})p_{1|t}(z_{1}|Z_{t}).$ (9)
We thus interpret the denoising as a two-step process: 1. Predict a clean instance from the noisy data, $p_{1|t}(z_{1}|Z_{t})$, 2. Re-noise the predicted clean instance, $q_{s|1}(z|z_{1})$. The main difference from the standard DDM model is that $z_{s}$ in ID directly depends only on the predicted clean instance, unlike the DDM model in which it directly depends on both the predicted clean instance and $z_{t}$. In Figure LABEL:fig:denoising we juxtapose the denoising graphical models of DDM and ID. The fact that $z_{s}$ is generated only from the predicted clean instance acts as a barrier to error propagation. This factorization will also prove useful in Section 4.
At inference, we define the Number of Function Evaluations (NFE), denoted as $T$, as a hyperparameter. The time step size is set to $\Delta_{T}=1/T$. The whole denoising process consists of iterating $T$ times the denoising step defined in Equation 8.
@@ 3.2.1 Corrector Sampling Interpretation
The denoising process can also be viewed through the lens of discrete diffusion corrector sampling. Defining a discrete diffusion corrector sampling step as applying $k\Delta_{t}$ backward steps, and $(k-1)\Delta_{t}$ forward steps, and noticing that the maximal corrector step is such that $11-k\Delta_{t}=$ 1, we state the following proposition:
@@
An Iterative Denoising step, as described in Equation 8, is equivalent to a discrete diffusion corrector sampling step with maximal corrector step size.
@@
Consequently, the denoising process inherits the properties of the backward process in discrete diffusion; by iteratively sampling $p_{s|t}(z|z_{t})$ with sufficiently small $\Delta_{t}$, the denoising process gradually transitions from the noise distribution to the data distribution.
@@ 3.3 Parametrization and Learning
Similar to discrete diffusion, the conditional probability $p_{1|t}(z_{1}|Z_{t})$ is intractable. We model this distribution by parametrizing a neural network $f_{\theta}(Z_{t},\alpha_{t})$ called the denoiser. We use Graph Neural Networks (GNNs) to enforce equivariance to node permutations, and thereby preserving the essential structural invariance of graphs.
Since the training objective is identical to that of DDM and DFM, we can adopt any training criterion proposed for these models to train $f_{\theta}(Z_{t},\alpha_{t})$. Similar to Digress (Vignac et al., 2023), we minimize the weighted negative log-likelihhood in our experiments:
${\mathcal{L}}=$ $\mathbb{E}_{{\mathcal{G}}\sim p_{\text{data}},t\sim{\mathcal{U}}(0,1),{\mathcal{G}}_{t}\sim(q_{t|1}(x_{t}^{(i)}|x^{(i)}_{1}),q_{t|1}(e^{(i,j)}_{t}|e^{(i,j)}_{1}))}$ (10)
[.$\biggl{[}\biggl{.}\gamma\sum_{x^{(i)}_{1}}\left[-\text{log}(p_{\theta}(x^{(i)}_{1}|{\mathcal{G}}_{t}))\right]$
$+(1-\gamma)\sum_{e^{(i,j)}_{1}}\left[-\text{log}(p_{\theta}(e^{(i,j)}_{1}|{\mathcal{G}}_{t}))\right]\biggr{.}\biggr{]},$
where $\gamma$ is a weighting factor between nodes and edges. We use $\gamma=n/(n+m)$. We can leverage any pre-trained DDM model for inference in the Iterative Denoising (ID) framework.
4 Critical Denoising
This section introduces Critical Iterative Denoising (CID), a method designed to improve Iterative Denoising.
As discussed in Section 3, Mask ID can be viewed as a sequence of denoising steps comprising: (1) unmasking elements of an instance based on the denoiser’s prediction and (2) re-masking a random subset of the instance’s elements.
However, in Mask ID, all elements are re-masked with the same probability. This approach is suboptimal, as not all elements are equally likely under the data distribution after the unmasking step.
CID addresses this limitation by dynamically adjusting the re-noising probability during sampling. Specifically, denoised elements that are overrepresented under the data distribution are re-noised with a higher probability, while elements that are underrepresented are more likely to remain uncorrupted. Our objective here is twofold: (1) Reducing error propagation by resampling elements that are likely out of the data distribution, and (2) Accelerating inference by decreasing the number of function evaluations required during sampling.
Figure 1 illustrates that CID achieves a state-of-the-art validity rate on Zinc250k with only a few tens of denoising steps.
@@ 4.1 Preliminaries
Before proceeding to the description of the method, we will recast the definition/computation of the noise distributions by introducing a Bernoulli random variable $a_{t}$ the mean of which is given by the $\alpha_{t}$ scheduling parameter in Equation 6; $a_{t}$ indicate whether an element has been corrupted ($a_{t}=0$) or not ($a_{t}=1)$. As with $z$, we use the capital letter $A$ to denote the set of indicators for all elements in an instance. The resulting formulation allows us to interpret the noising noising process as selectively corrupting some elements while leaving others untouched. Through the random variables $a_{t}$, we can identify which are the $z_{t}$ elements that have been noised. This identification would otherwise be impossible for any noise distribution sharing support with the data distribution (i.e., all practical noise distributions except the mask distribution).
Thanks to the random variable $a_{t}\sim p_{\alpha_{t}}(a)=\text{Bernoulli}(a;\alpha_{t})$, we rewrite Equation 6 as:
$q_{t}(z|z_{1})=p_{\alpha_{t}}(a)\delta_{z_{1}}(z_{t})+(1-p_{\alpha_{t}}(a))q_{0}(z).$ (11)
Similarly, we rewrite the denoising equation (Eq. 8) as:
$p_{s|t}(z|Z_{t})=p_{\alpha_{t}}(a)\delta_{z_{t}}(z)+(1-p_{\alpha_{t}}(a))p^{\theta}_{1|t}(z|Z_{t},A_{t})$ (12)
Intuitively, when the element is not corrupted ($a_{t}=1$,) it remains unchanged ($\delta_{z_{t}}(z)$); when $a_{t}=0$ we denoise it using the denoising distribution $p^{\theta}_{1|t}(z|Z_{t},A_{t})$. We simplify notation as follows $p_{\text{data}}(z):=\delta_{z_{t}}(z)$ and $p_{\text{pred}}(z):=p^{\theta}_{1|t}(z|Z_{t},A_{t})$.
In principle, our method supports any noise distribution and produces pairs $(z_{t},a_{t})$, which indicate whether the element $z_{t}$ was sampled from the noise distribution or corresponds to the original clean element. However, denoising using the pair $(z_{t},a_{t})$ reduces denoising models with with any noise distribution (e.g., marginal, uniform) to a mask denoising model.
To see this, we observe that the corruption indicator $a_{t}$ acts as a mask, with $a_{t}=0$ indicating a masked element. Defining the set of uncorrupted elements as $Z^{a}_{t}=\{z_{t}\in Z_{t}\mid a_{t}=1\}$, and its complement as $Z^{\bar{a}}_{t}$, the masked denoiser can be expressed as $p^{\theta}_{1\mid t}(z\mid Z^{a}_{t},A_{t})$. Since $Z^{\bar{a}}_{t}$ carries no information about $Z_{1}$, it follows that, for any noise distribution, $p^{\theta}_{1\mid t}(z\mid Z_{t},A_{t})=p^{\theta}_{1\mid t}(z\mid Z^{a}_{t},A_{t}).$ As Critical Iterative Denoising (CID) explicitly requires $A_{t}$, and any noise distribution under this configuration collapses to the Mask Iterative Denoising (Mask ID) model, we always use Mask ID within CID.
@@ 4.2 Iterative Denoising with a Critic
We now describe the critic used to predict the corruption state of the graph elements during the denoising phase. Let $\hat{z}_{1|t}\sim p_{1|t}(z|Z_{t},A_{t})$ denote an element after the denoising step at time $t$. CID trains a Critic $C$ that models $\hat{\alpha}_{t}=p_{\phi}(a|\hat{Z}_{1|t})$, which predicts the probability that $\hat{z}$ comes from $p_{\text{data}}$ or $p_{\text{pred}}$. The Critic is trained to minimize the negative log-likelihood of the $a^{(i)}_{t}$ corruption indicators produced at the noising phase (one per element and time step):
${\mathcal{L}}_{\phi}=-\mathbb{E}_{t,\hat{Z}_{1|t}}\sum_{i}\log p_{\phi}(a^{(i)}_{t}|\hat{Z}_{1|t}),$ (13)
During inference, $\hat{\alpha}_{t}$ parametrizes the Bernoulli distribution $p_{\hat{\alpha}_{t}}(a_{t})=\text{Bernoulli}(a;\hat{\alpha}_{t})$ which determines the noising probability $1-\hat{\alpha}_{t}$. It is important to remind that $\alpha_{t}$ is the sheduler’s noise rate at time step $t$. So we interpret the output of the critic as an adaptive scheduling parameter setting the level of noise per element.
In the following theorem we characterise the optimal critic:
@@
$p_{data}(\hat{z}_{1|t})=p_{pred}(\hat{z}_{1|t})\implies\hat{\alpha}^{*}_{t}=\alpha_{t}$ (15)
@@
If $p_{data}(\hat{z}_{1|t})>p_{pred}(\hat{z}_{1|t})$, the optimal Critic’s noising probability $\hat{\alpha}_{t}^{*}$ is smaller than the schedule’s noising rate $\alpha_{t}$, that is:
$p_{data}(\hat{z}_{1|t})>p_{pred}(\hat{z}_{1|t})\implies\alpha_{t}>\hat{\alpha}_{t}^{*}.$ (16)
Conversely, if $p_{data}(\hat{z}_{1|t})<p_{pred}(\hat{z}_{1|t})$, the optimal Critic’s noising probability $\hat{\alpha}_{t}^{*}$ is larger than the schedule’s noising rate $\alpha_{t}$:
$p_{data}(\hat{z}_{1|t})<p_{pred}(\hat{z}_{1|t})\implies\alpha_{t}<\hat{\alpha}_{t}^{*}.$ (17)
@@
Additionally, if $p_{\text{pred}}(\hat{z}_{1|t})=0$ the element will not be masked ($\alpha_{t}=0$). Conversely, if $p_{\text{data}}(\hat{z}_{1|t})=0$, the element will be masked with probability $\alpha_{t}=1$.
Thanks to Theorem 4.1 and the following lemmas, the Critic steers the denoised element toward the data distribution, by re-noising with higher probability elements that are overrepresented under the data distribution, and leaving uncorrupted elements that are underrepresented.

### v2

@@ 3.1 Noising
$q_{t|1}(z\mid z_{1})=\alpha_{t}\delta_{z_{1}}(z)+(1-\alpha_{t})q_{0}(z),$ (6)
where $q_{0}(z)$ represents the noise distribution $\delta_{z_{1}}(z)$ is the data distribution, and $\alpha_{t}$ is a non-decreasing scheduling parameter in $[0,1]$. The scheduler $\alpha_{t}$, which defines the noise level, depends on a continuous time parameter, which also takes values in $[0,1]$.
Unlike standard diffusion models, which compute noise distributions via a Markov process (Eq. 1) such that $q_{t|t+\Delta_{t},1}(z\mid z_{t+\Delta_{t}},z_{1})=q_{t|t+\Delta_{t}}(z\mid z_{t+\Delta_{t}})$, our intermediate noisy distributions are directly parameterized as a mixture of the noise and the data distribution. Thus, we do not assume any dependencies in the noising process other than the dependency on the original data point $z_{1}$ and on the noise level. This assumption introduces a significant simplification over DDM and DFM, leading to beneficial implications for the SID denoising process. In Figure LABEL:fig:noising, we juxtapose the graphical models of our method and DDM at the noising phase. More formally we assume conditional independence of the noisy distributions across time given $z_{1}$:
@@
The intermediate noisy distribution at time $t$ is conditionally independent from other intermediate states such that:
$q_{t|1}(z\mid z_{1})=q_{t|1}(z\mid z_{s},z_{1})\quad\forall\>t\neq s.$ (7)
As a result, our noising procedure is not a diffusion process in the traditional sense, as the state at time $t$ depends only on the initial state and not on the noising trajectory or previous intermediate states. Under this assumption, the denoising phase does not suffer from error propagation and accumulation caused by the direct dependencies to previous states inherent in standard DDM and DFM (see Section 3.2).
@@
Given a noise distribution $q_{0}$ and a scheduler $\alpha(t)$, the distribution $q^{SID}_{t|1}(z\mid z_{1})$ defined by our noising procedure (Equation 6) is equal to the distribution $q^{DDM}_{t|1}(z\mid z_{1})$ obtained by the discrete forward process distribution defined in Equation 1.
@@
The proposition follows from the fact that our noising procedure corresponds to the closed-form expression (Eq. 3) used in discrete diffusion to sample efficiently from $q_{t|1}(z\mid z_{1})$. However, we make no assumption that the noisy distribution results from a diffusion process.
@@ 3.2 Denoising
Denoising progressively refines noisy instances by learning the distribution $p_{s|t}(z\mid Z_{t})$, where $s=t+\Delta t$. By leveraging Assumption 3.1, the denoising process is significantly simplified compared to discrete diffusion, as formalized in the following proposition:
@@
Under Assumption 3.1, the denoising process factorizes as:
$p_{s|t}(z\mid Z_{t})=\sum_{z_{1}}p_{1|t}(z_{1}\mid Z_{t})q_{s|1}(z\mid z_{1}).$ (9)
Thus, we interpret the denoising as a two-step process: 1. Predict a clean instance from the noisy data, $p_{1|t}(z_{1}\mid Z_{t})$, 2. Re-noise the predicted clean instance, $q_{s|1}(z\mid z_{1})$. The key difference from the standard DDM model is that $z_{s}$ in SID directly depends only on the predicted clean instance, but not on $z_{t}$. In Figure LABEL:fig:denoising we juxtapose the denoising graphical models of DDM and SID. The fact that $z_{s}$ is generated only from the predicted clean instance acts as a barrier to error propagation. This factorization will also prove useful in Section 4.
At inference time, we define the Number of Function Evaluations (NFE), denoted as $T$, as a hyperparameter. The time step size is set to $\Delta_{t}=1/T$. The whole denoising process consists of iterating $T$ times the denoising step defined in Equation 8.
@@ 3.2.1 Corrector Sampling Interpretation
The denoising process can also be viewed through the lens of discrete diffusion corrector sampling. Defining a discrete diffusion corrector sampling step as applying $k\Delta_{t}$ backward steps and $(k-1)\Delta_{t}$ forward steps, and noticing that the maximal size for corrector step at time $t$ is $k\Delta_{t}=1-t$, we state the following proposition:
@@
A Simple Iterative Denoising step, as described in Equation 8, is equivalent to a discrete diffusion corrector sampling with maximal corrector step sizes.
@@
Consequently, the denoising process inherits the properties of the backward process in discrete diffusion; by iteratively sampling $p_{s|t}(z\mid Z_{t})$ with sufficiently small $\Delta_{t}$, the denoising process gradually transitions from the noise distribution to the data distribution. Our mask SID also inherits the time-independence property of mask denoisers shown by Ou et al. (2025) and Sahoo et al. (2024).
@@ 3.3 Parametrization and Learning
As in discrete diffusion, the conditional probability $p_{1|t}(z_{1}\mid Z_{t})$ is intractable. We model this distribution by parametrizing a neural network $f_{\theta}(Z_{t},\alpha_{t})$ referred to as the denoiser. We use Graph Neural Networks (GNNs) to enforce equivariance to node permutations, thereby preserving the structural invariance of graphs.
Since the training objective matches that of DDM and DFM, we can adopt any of the training criteria proposed for these models to train $f_{\theta}(Z_{t},\alpha_{t})$. Following Digress (Vignac et al., 2023), we minimize the weighted negative log-likelihood in our experiments:
${\mathcal{L}}=$ $\mathbb{E}_{{\mathcal{G}}\sim p_{\text{data}},t\sim{\mathcal{U}}(0,1),{\mathcal{G}}_{t}\sim(q_{t|1}(x_{t}^{(i)}\mid x^{(i)}_{1}),q_{t|1}(e^{(i,j)}_{t}\mid e^{(i,j)}_{1}))}$ (10)
[.$\biggl{[}\biggl{.}\gamma\sum_{x^{(i)}_{1}}\left[-\log(p_{\theta}(x^{(i)}_{1}\mid{\mathcal{G}}_{t}))\right]$
$+(1-\gamma)\sum_{e^{(i,j)}_{1}}\left[-\log(p_{\theta}(e^{(i,j)}_{1}|{\mathcal{G}}_{t}))\right]\biggr{.}\biggr{]},$
where $\gamma$ is a weighting factor between nodes and edges. We use $\gamma=n/(n+m)$. At inference time, Simple Iterative Denoising can leverage any pre-trained DDM model.
4 Critical Iterative Denoising
This section introduces Critical Iterative Denoising (CID), a method designed to improve Simple Iterative Denoising.
As discussed in Section 3, Mask SID can be viewed as a sequence of denoising steps comprising: (1) unmasking elements of an instance based on the denoiser’s prediction and (2) re-masking a random subset of the instance’s elements.
However, in Mask SID, all elements are re-masked with the same probability. This approach is suboptimal, as not all elements are equally likely under the data distribution after the unmasking step.
CID addresses this limitation by dynamically adjusting the re-noising probability during sampling. Specifically, CID increases the re-noising probability for predicted denoised elements overrepresented under the data distribution and decreases it for those underrepresented. CID serves two goals: (1) Reducing error propagation by resampling elements that are likely out of the data distribution, and (2) Accelerating inference by decreasing the number of function evaluations required during sampling. As shown in Figure 1, CID achieves a state-of-the-art validity rate on Zinc250k with only a few dozen of denoising steps.
@@ 4.1 Preliminaries
Before describing the method, we recast the definition of the noise distributions by introducing a Bernoulli random variable $a_{t}$, whose mean is given by the scheduling parameter $\alpha_{t}$ from Equation 6; $a_{t}$ indicates whether an element has been corrupted ($a_{t}=0$) or not ($a_{t}=1)$. As with $z$, we use the capital letter $A$ to denote the set of indicators for all elements in an instance. This formulation allows us to interpret the noising process as selectively corrupting some elements while leaving others untouched. Through the random variables $a_{t}$, we can identify which elements $z_{t}$ have been noised. This identification would otherwise be impossible for any noise distribution sharing support with the data distribution (i.e., all practical noise distributions except the mask distribution).
Given the random variable $a_{t}\sim p_{\alpha_{t}}(a)=\text{Bernoulli}(a;\alpha_{t})$, we rewrite Equation 6 as:
$q_{t|1}(z|z_{1})=p_{\alpha_{t}}(a)\delta_{z_{1}}(z)+(1-p_{\alpha_{t}}(a))q_{0}(z).$ (11)
Similarly, we write the denoising equation as:
$p_{1|t}(z\mid Z_{t},A_{t})=a_{t}\delta_{z_{1}}(z)+(1-a_{t})p^{\theta}_{1|t}(z\mid Z_{t},A_{t})$ (12)
Intuitively, when the element is not corrupted ($a_{t}=1$), it remains unchanged ($\delta_{z_{1}}(z)$); when $a_{t}=0$ we predict it using the denoiser. For simplicity, we define $p_{\text{data}}(z):=\delta_{z_{1}}(z)$ and $p_{\text{pred}}(z):=p^{\theta}_{1|t}(z\mid Z_{t},A_{t})$.
In principle, our method supports any noise distribution and produces pairs $(z_{t},a_{t})$, which indicate whether the element $z_{t}$ was sampled from the noise distribution or corresponds to the original clean element. However, denoising using the pair $(z_{t},a_{t})$ reduces denoising models with any noise distribution (e.g., marginal, uniform) to a mask denoising model.
To see this, note that the corruption indicator $a_{t}$ acts as a mask, with $a_{t}=0$ indicating a masked element. Defining the set of uncorrupted elements as $Z^{a}_{t}=\{z_{t}\in Z_{t}\mid a_{t}=1\}$, and its complement as $Z^{\bar{a}}_{t}$, the masked denoiser can be expressed as $p^{\theta}_{1\mid t}(z\mid Z^{a}_{t},A_{t})$. Since $Z^{\bar{a}}_{t}$ carries no information about $Z_{1}$, it follows that: $p^{\theta}_{1\mid t}(z\mid Z_{t},A_{t})=p^{\theta}_{1\mid t}(z\mid Z^{a}_{t},A_{t}).$ Because Critical Iterative Denoising (CID) explicitly requires $A_{t}$, and any noise distribution under this configuration collapses to the Mask Simple Iterative Denoising (Mask SID) model, we always use Mask SID within CID.
@@ 4.2 Iterative Denoising with a Critic
We now describe the Critic, which predicts the corruption state of graph elements during denoising. Let $\hat{z}_{1|t}\sim p_{1|t}(z\mid Z_{t},A_{t})$ denote a denoised element at time $t$. CID trains a Critic $C$ to estimate $\hat{\alpha}_{t}=p_{\phi}(a\mid\hat{Z}_{1|t})$, the probability that $\hat{z}$ originates from $p_{\text{data}}$ rather than $p_{\text{pred}}$. The Critic is trained by minimizing the negative log-likelihood of the corruption indicators $a^{(i)}_{t}$ produced during the noising phase (one per element and time step):
${\mathcal{L}}_{\phi}=-\mathbb{E}_{t,\hat{Z}_{1|t}}\sum_{i}\log p_{\phi}(a^{(i)}_{t}\mid\hat{Z}_{1|t}),$ (13)
During inference, $\hat{\alpha}_{t}$ parametrizes a Bernoulli distribution $p_{\hat{\alpha}_{t}}(a_{t})=\text{Bernoulli}(a;\hat{\alpha}_{t})$ which determines the noising probability $1-\hat{\alpha}_{t}$. Recall that $\alpha_{t}$ is the scheduler’s noise rate at time step $t$. Thus, the Critic’s output can be interpreted as an adaptive, element-wise noise schedule.
The following theorem characterizes the optimal critic:
@@
$p_{\text{data}}(\hat{z}_{1|t})=p_{\text{pred}}(\hat{z}_{1|t})\implies\hat{\alpha}^{*}_{t}=\alpha_{t}$ (15)
@@
If $p_{\text{data}}(\hat{z}_{1|t})>p_{\text{pred}}(\hat{z}_{1|t})$, the optimal Critic’s noising probability $\hat{\alpha}_{t}^{*}$ is smaller than the schedule’s noising rate $\alpha_{t}$, that is:
$p_{\text{data}}(\hat{z}_{1|t})>p_{\text{pred}}(\hat{z}_{1|t})\implies\alpha_{t}>\hat{\alpha}_{t}^{*}.$ (16)
Conversely, if $p_{\text{data}}(\hat{z}_{1|t})<p_{\text{pred}}(\hat{z}_{1|t})$, the optimal Critic’s noising probability $\hat{\alpha}_{t}^{*}$ is larger than the schedule’s noising rate $\alpha_{t}$:
$p_{\text{data}}(\hat{z}_{1|t})<p_{\text{pred}}(\hat{z}_{1|t})\implies\alpha_{t}<\hat{\alpha}_{t}^{*}.$ (17)
@@
Additionally, if $p_{\text{pred}}(\hat{z}_{1|t})=0$, then $\hat{z}_{1|t}$ will not be masked ($\alpha_{t}=0$). Conversely, if $p_{\text{data}}(\hat{z}_{1|t})=0$, the element is out of distribution and will be masked with probability $\alpha_{t}=1$.
Thanks to Theorem 4.1 and the following lemmas, the Critic steers the denoised element toward the data distribution, by re-noising with higher probability those elements that are overrepresented under the data distribution, while preserving underrepresented ones.